OpenStudy (ijlal):

Hello :) I am having trouble integrating this problem it's related with a part of Gamma functions and Fresnel Integral

OpenStudy (ijlal):

\[\int\limits_{0}^{\infty}\sin(u^2)du\]

OpenStudy (ijlal):

@Photon336 @mathstudent55

OpenStudy (ijlal):

is there a specific gamma function formula for these type of integrals?

OpenStudy (inkyvoyd):

" The limits of these functions as x goes to infinity are known:..."

OpenStudy (inkyvoyd):

Wikipedia can offer you the answer but not the process I guess...

OpenStudy (inkyvoyd):

There are some formulas for the integrals in terms of erf and erfi IIRC

OpenStudy (ijlal):

@inkyvoyd thank you for sharing this link i looked it up it's in the definition of error functions and elementary functions but i need to know how do we integrate it like what are the steps to it using gamma functions

OpenStudy (ijlal):

Please anyone some help would be appreciated. You also provide links of videos to help me understand i tried and search most of them were too complicated and direct

OpenStudy (ijlal):

@Photon336 hey buddy do you know anyone who could help me with this :/

OpenStudy (ijlal):

like is there any direct formula of gamma functions on how to evaluate these type of integrals? or is there a process as in steps?

OpenStudy (irishboy123):

totally talking out of my as5 here but can you do this as the imaginary part of the full integration and it's a definite integral so you can Feynman it too. i know that sounds vague, lol, but if i actually knew how to do it, i would tell you!!!

OpenStudy (ijlal):

i remember a bit but forgot \[\frac{ \Gamma(s) }{ n|p|^s }\sin (\alpha s)= ?\]

OpenStudy (ijlal):

i forgot what the right hand side is equal to can anyone tell me?

OpenStudy (irishboy123):

wow!! that looks awesome!!

OpenStudy (photon336):

@Kainui

OpenStudy (ijlal):

@IrishBoy123 haha :P i still remember when maths used to be just numbers the good old days :')

OpenStudy (ijlal):

and how do we find the value of s p and n on the left hand side ? any1?

OpenStudy (irishboy123):

lol!!! i am going to grab my Abacus and then we can talk business!!!!

OpenStudy (ijlal):

hahah xD

OpenStudy (ijlal):

never mind found it out :P thanks though for all your help :) here is the video that explains this funny integral if you like to see :D https://www.youtube.com/watch?v=H0kT-EKbUzM